### v1

@@ 3.3.2 Multimodal Sentiment Semantic Fusion
Tensor-based fusion. In the early stages of audio-image-text sentiment analysis research, considering the small scale of datasets and limited computational resources, researchers represented the raw inputs of each modality as a single emotional semantic representation before proceeding to multimodal emotional semantic representation fusion. The simplest fusion strategy was to directly concatenate the emotional semantic representations of different modalities, but this method did not explicitly model the higher-order interactions between emotional information from different modalities. To address this issue, in [33], the authors proposed using the outer product of vectors to fuse different modal representations, thereby modeling interactions among unimodal, bimodal, and trimodal emotional semantic representations simultaneously. However, this method, due to the complexity of the outer product operation being tied to the product of input vector dimensions, resulted in high computational costs and slow efficiency. Subsequent work has made efficiency improvements. In [187], the authors proposed the LMF fusion method, which accelerates the fusion process of multimodal emotional representations through low-rank decomposition. In [188], the authors introduced a three-stage multimodal emotional representation fusion strategy consisting of representation slice grouping, intra-group representation slice fusion, and global representation fusion. Representation slice grouping involves splitting the representations of each modality into the same fixed number of small groups, numbering them, and then locally fusing representation slices of the same number from different modalities together. This approach reduces the dimensions of representations to be fused later, thereby enhancing fusion efficiency. Intra-group representation slice fusion uses the outer product method to fuse the representation slices of the three modalities within the group, which, due to the smaller feature dimensions, significantly speeds up the fusion process. Finally, Long Short-Term Memory (LSTM) networks are used to perform global representation fusion of the different groups after fusion. This method reduces the computational complexity of the tensor outer product fusion method to some extent through block processing.
@@
Fine-grained temporal interaction modeling fusion. This type of fusion method focuses on capturing more localized, fine-grained interactions of multimodal information. These methods first obtain fine-grained representations corresponding to each time step of each modality, and then perform multimodal sentiment semantic fusion based on these representations to capture the interactions between cross-modal and cross-temporal sentiment information. In [34], the RAVEN model is a typical method in this series of research. The authors found that the same words can convey different emotional messages when accompanied by different tones or expressions. Driven by this motivation, they designed a network that improves the word representations by dynamically integrating the fine-grained representations of visual and auditory modalities into each word vector through a cross-modal gating mechanism, thereby achieving the goal of infusing non-verbal emotional information into word representations. In[189], considering that audio and visual inputs might contain noise at certain time steps, like background noise in speech, the authors proposed a reinforcement learning-based gating unit to control the information fusion between fine-grained representations of different modalities. The gating mechanism allows for dynamic sentiment representation fusion by controlling whether the representation of the current word incorporates information from a particular modality. Unlike the previous two works, which focus on capturing interactions of multimodal fine-grained sentiment representations associated with individual words, in [190], the authors model the feature interactions between multimodal fine-grained sentiment representations of multiple consecutive words within a window and use a memory neural network to model global information.
@@
To effectively extend LLMs and LMMs to multimodal sentiment analysis tasks and address two pressing challenges in the field: (1) the low contribution rate of the visual modality and (2) the design of an effective multimodal fusion architecture, scholars [204] have proposed an inter-frame hybrid transformer. This transformer extracts spatiotemporal features from sparsely sampled video frames, focusing not only on facial expressions but also capturing body movement information.
@@ 3.4 Multimodal Sarcasm Detection
Additionally, some researchers have studied sarcasm in text and images; for example, in [55], the authors introduced a multimodal sarcasm detection task for text and images and designed a multi-level fusion network to detect sarcasm. In [194], the authors proposed the multimodal sarcasm detection model using two different computational frameworks based on SVM and CNN that integrate text and visual modalities. Identifying inconsistencies between modalities is key to multimodal sarcasm detection, and recent models can be categorized into those based on attention mechanisms and those using Graph Neural Networks (GNN). For example, in [195], the authors introduced a BERT-based model with a cross-modal attention mechanism and a text-oriented co-attention mechanism to capture inconsistencies within and between modalities. In [196], the authors designed a 2D internal attention mechanism based on BERT and ResNet to extract relationships between words and images. In [197], the authors proposed a Transformer-based architecture to fuse textual and visual information. In terms of GNN, scholars in [198] built heterogeneous intramodal and cross-modal graphs (InCrossMG) for each multimodal example to determine the emotional inconsistencies within specific modalities and between different modalities, and introduced an interactive graph convolutional network structure to learn the relationships of inconsistencies in a joint and interactive manner within modal and cross-modal graphs. In [199], the authors constructed heterogeneous graphs containing fine-grained object information of images for each instance and designed a cross-modal graph convolutional network. Additionally, some scholars [200] have proposed incorporating prior knowledge into multimodal sarcasm detection. They introduced KnowleNet, which utilizes the ConceptNet [201] knowledge base to integrate prior knowledge, and determines the relevance between images and text through sample-level and word-level cross-modal semantic similarity detection. They also incorporated contrastive learning to improve the spatial distribution of sarcastic (positive) and non-sarcastic (negative) samples. In [202], the authors propose a lightweight multimodal interaction model with knowledge enhancement based on deep learning. In [203], the authors proposed to introduce emotional knowledge into sarcasm detection and used sentiment dictionaries to obtain the sentiment vectors by evaluating the words extracted from various modalities, and then combined them with each modality. In [205], the authors propsed multi-view CLIP that is capable of leveraging multi-grained cues from multiple perspectives (i.e., text, image, and textimage interaction view) for multi-modal sarcasm detection.
@@ 3.5 The Usage of LLMs in Multimodal Sentiment Analysis
4 LLMs and LMMs in Multimodal Sentiment Analysis Evaluations
@@ 4.3 Reference Results
With the in-depth development of LLMs in the field of multimodal sentiment analysis, it is necessary to compare the performance of LLMs on multimodal sentiment analysis datasets. However, testing on commercial LLMs such as ChatGPT is often expensive. Some works [141, 160, 171, 206, 207, 213, 214] have demonstrated the performance of some LLMs on multimodal sentiment analysis tasks.

### v2

@@ 3.3.2 Multimodal Sentiment Semantic Fusion
Tensor-based fusion. In the early stages of audio-image-text sentiment analysis research, considering the small scale of datasets and limited computational resources, researchers represented the raw inputs of each modality as a single emotional semantic representation before proceeding to multimodal emotional semantic representation fusion. The simplest fusion strategy was to directly concatenate the emotional semantic representations of different modalities, but this method did not explicitly model the higher-order interactions between emotional information from different modalities. To address this issue, in [33], the authors proposed using the outer product of vectors to fuse different modal representations, thereby modeling interactions among unimodal, bimodal, and trimodal emotional semantic representations simultaneously. However, this method, due to the complexity of the outer product operation being tied to the product of input vector dimensions, resulted in high computational costs and slow efficiency. Subsequent work has made efficiency improvements. In [187], the authors proposed the LMF fusion method, which accelerates the fusion process of multimodal emotional representations through low-rank decomposition.
Figure 8: Decomposing weight tensor into low-rank factors [187].
In [188], the authors introduced a three-stage multimodal emotional representation fusion strategy consisting of representation slice grouping, intra-group representation slice fusion, and global representation fusion. Representation slice grouping involves splitting the representations of each modality into the same fixed number of small groups, numbering them, and then locally fusing representation slices of the same number from different modalities together. This approach reduces the dimensions of representations to be fused later, thereby enhancing fusion efficiency. Intra-group representation slice fusion uses the outer product method to fuse the representation slices of the three modalities within the group, which, due to the smaller feature dimensions, significantly speeds up the fusion process. Finally, Long Short-Term Memory (LSTM) networks are used to perform global representation fusion of the different groups after fusion. This method reduces the computational complexity of the tensor outer product fusion method to some extent through block processing.
@@
Fine-grained temporal interaction modeling fusion. This type of fusion method focuses on capturing more localized, fine-grained interactions of multimodal information. These methods first obtain fine-grained representations corresponding to each time step of each modality, and then perform multimodal sentiment semantic fusion based on these representations to capture the interactions between cross-modal and cross-temporal sentiment information. In [34], the RAVEN model is a typical method in this series of research. The authors found that the same words can convey different emotional messages when accompanied by different tones or expressions. Driven by this motivation, they designed a network that improves the word representations by dynamically integrating the fine-grained representations of visual and auditory modalities into each word vector through a cross-modal gating mechanism, thereby achieving the goal of infusing non-verbal emotional information into word representations. For a word $\mathbf{L}^{(i)}$, the nonverbal shift vector $\mathbf{h}^{(i)}_{m}$ is calculated as follows:
$\mathbf{h}^{(i)}_{m}=w_{v}^{(i)}\cdot(\mathbf{W}_{v}\mathbf{h}^{(i)}_{v})+w^{(i)}_{a}\cdot(\mathbf{W}_{a}\mathbf{h}^{(i)}_{a})+\mathbf{b}^{(i)}_{h}$ (18)
where $\mathbf{W}_{v}$ and $\mathbf{W}_{a}$ are weight matrices for the visual and acoustic embedding and $\mathbf{b}^{(i)}_{h}$ is the bias vector.
In[189], considering that audio and visual inputs might contain noise at certain time steps, like background noise in speech, the authors proposed a reinforcement learning-based gating unit to control the information fusion between fine-grained representations of different modalities. The gating mechanism allows for dynamic sentiment representation fusion by controlling whether the representation of the current word incorporates information from a particular modality. Unlike the previous two works, which focus on capturing interactions of multimodal fine-grained sentiment representations associated with individual words, in [190], the authors model the feature interactions between multimodal fine-grained sentiment representations of multiple consecutive words within a window and use a memory neural network to model global information.
@@
Figure 9: Overview of the VLP2MSA [204] architecture(a) and the Inter-frame Hybrid Transformer module(b). The model receives the original text and video frames as input and consists of a text encoder, a video encoder, a prompt module, a video-text contrastive, and a multimodal integration encoder. we propose an adapted and extended video encoder and a new fusion scheme that prompts and aligns before interacting to reduce the influence of modality heterogeneity on the fusion results.
To effectively extend LLMs and LMMs to multimodal sentiment analysis tasks and address two pressing challenges in the field: (1) the low contribution rate of the visual modality and (2) the design of an effective multimodal fusion architecture, scholars [204] have proposed an inter-frame hybrid transformer. As shown in Figure 9, this transformer extracts spatiotemporal features from sparsely sampled video frames, focusing not only on facial expressions but also capturing body movement information.
@@ 3.4 Multimodal Sarcasm Detection
Additionally, some researchers have studied sarcasm in text and images; for example, in [55], the authors introduced a multimodal sarcasm detection task for text and images and designed a multi-level fusion network to detect sarcasm. In [194], the authors proposed the multimodal sarcasm detection model using two different computational frameworks based on SVM and CNN that integrate text and visual modalities. Identifying inconsistencies between modalities is key to multimodal sarcasm detection, and recent models can be categorized into those based on attention mechanisms and those using Graph Neural Networks (GNN). For example, in [195], the authors introduced a BERT-based model with a cross-modal attention mechanism and a text-oriented co-attention mechanism to capture inconsistencies within and between modalities. In [196], the authors designed a 2D internal attention mechanism based on BERT and ResNet to extract relationships between words and images. In [197], the authors proposed a Transformer-based architecture to fuse textual and visual information. In terms of GNN, scholars in [198] built heterogeneous intramodal and cross-modal graphs (InCrossMG) for each multimodal example to determine the emotional inconsistencies within specific modalities and between different modalities, and introduced an interactive graph convolutional network structure to learn the relationships of inconsistencies in a joint and interactive manner within modal and cross-modal graphs. In [199], the authors constructed heterogeneous graphs containing fine-grained object information of images for each instance and designed a cross-modal graph convolutional network.
Figure 10: Different processing procedures for positive(sarcasm) and negative(non-sarcasm) samples in knowledge-based word-level cross-modal semantic similarity detection module [200].
Additionally, some scholars [200] have proposed incorporating prior knowledge into multimodal sarcasm detection. They introduced KnowleNet, which utilizes the ConceptNet [201] knowledge base to integrate prior knowledge, and determines the relevance between images and text through sample-level and word-level cross-modal semantic similarity detection. They also incorporated contrastive learning to improve the spatial distribution of sarcastic (positive) and non-sarcastic (negative) samples. In [202], the authors propose a lightweight multimodal interaction model with knowledge enhancement based on deep learning. In [203], the authors proposed to introduce emotional knowledge into sarcasm detection and used sentiment dictionaries to obtain the sentiment vectors by evaluating the words extracted from various modalities, and then combined them with each modality. In [205], the authors propsed multi-view CLIP that is capable of leveraging multi-grained cues from multiple perspectives (i.e., text, image, and textimage interaction view) for multi-modal sarcasm detection.
Figure 11: An overview of CofiPara for multimodal sarcasm target identification [207].
@@ 3.5 The Usage of LLMs in Multimodal Sentiment Analysis
4 Evaluations of LLMs-based Multimodal Sentiment Analysis Methods
@@ 4.3 Reference Results
With the in-depth development of LLMs in the field of multimodal sentiment analysis, it is necessary to compare the performance of LLMs on multimodal sentiment analysis datasets. However, testing on commercial LLMs such as ChatGPT is often expensive. Some works [141, 160, 171, 206, 207, 213, 214, 217] have demonstrated the performance of some LLMs on multimodal sentiment analysis tasks, We have organized the relevant results in the table 4.
Table 4: An overview of the performance of existing LLMs and LMMs on text-centric multimodal sentiment analysis benchmarks. * indicates the model results after training on the MMSD and MMSD2.0 dataset. Italicized words represent the few-shot results, and the rest are zero-shot results. Bold represents the best zero-shot results.
Method MVSA-S MVSA-M TumEmo Twitter-2015 Twitter-2017 MMSD MMSD2.0 MOSI-2 MOSEI CH-SIMS M3ED
Acc. Acc. Acc. Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1
ChatGPT 56.55 53.18 48.17 65.28 63.12 62.47 62.34 69.02 - - - 86.13 85.92 85.60 84.43 79.66 78.78 44.47 40.40
Claude - - - - - - - - - - - 87.04 86.55 85.83 84.81 88.70 87.44 34.90 34.83
LLaMA1-7B 67.23 60.72 38.26 58.53 - 46.43 - 58.99 - - - 82.01 - 75.62 - - - - -
LLaMA1-13B 66.88 68.82 44.68 52.07 - 47.24 - 57.53 - - - 72.10 - 79.55 - - - - -
LLaMA2-7B 66.99 69.22 40.28 58.53 - 46.60 - 56.33 - - - 67.68 - 77.30 - - - - -
LLaMA2-13B 66.02 68.69 45.78 60.37 - 48.54 - 60.23 - - - 81.86 - 81.66 - - - - -
Mixtral-AWQ 54.37 55.59 43.94 55.45 - 60.21 - 64.38 - - - 87.92 - 79.30 - - - - -
Gemma 67.72 61.61 43.10 54.29 - 52.43 - 60.07 - - - 81.65 - 77.05 - - - - -
Flan-T5-XXL 64.81 66.01 49.56 72.13 - 63.70 - 71.40 - - - 89.60 - 86.52 - - - - -
ChatGLM2-6B - - - - - - - - - - - 84.12 84.12 - - 77.58 75.95 45.68 30.52
ChatGLM2-6B* - - - - - - - 94.02 93.76 78.41 78.23 - - - - - - - -
LLaMA2-7B* - - - - - - - 93.97 93.72 82.52 82.27 - - - - - - - -
GPT-4V 76.19 71.05 50.58 53.85 - 60.16 - 76.76 - - - 90.91 - 87.10 - - - - -
Claude3-V 80.95 69.08 46.15 38.46 - 54.47 - 71.37 - - - 78.79 - 79.93 - - - - -
Gemini-V 72.73 70.18 51.65 54.51 - 59.32 - 56.83 - - - 88.34 - 87.14 - - - - -
OpenFlamingo 55.58 61.15 29.47 57.28 - 46.19 - 52.68 - - - 79.97 - 77.30 - - - - -
Fromage 29.85 28.19 22.76 19.96 - 27.31 - 40.68 - - - 57.19 - 47.41 - - - - -
LLaVA-v0-7B 69.42 65.42 30.44 35.10 - 44.57 - 43.21 - - - 74.69 - 74.65 - - - - -
LLaVA-v0-13B 73.06 69.61 38.51 37.99 - 48.46 - 44.29 - - - 80.18 - 76.58 - - - - -
LLaVA-v1.6-7B 59.95 67.23 45.12 59.31 - 52.84 - 59.61 - - - 85.63 - 81.62 - - - - -
LLaVA-v1.6-13B 64.56 60.43 53.22 58.73 - 56.08 - 62.31 - - - 86.39 - 78.26 - - - - -
MiniGPT4 71.12 70.78 50.29 47.16 - 49.43 - 57.49 - - - 83.99 - 83.38 - - - - -
mPLOG-Owl 51.94 50.36 33.37 33.75 - 38.74 - 49.73 - - - 68.75 - 58.10 - - - - -
mPLOG-Owl2.1 53.64 63.11 47.02 60.66 - 55.11 - 60.48 - - - 85.63 - 73.47 - - - - -
AdapterV2 73.54 70.13 39.14 37.32 - 48.38 - 57.20 - - - 86.43 - 82.02 - - - - -
VPGTrans 64.32 69.54 46.17 42.62 - 44.81 - 65.04 - - - 76.22 - 76.76 - - - - -
MultiGPT 52.91 62.03 30.26 58.53 - 46.35 - 59.82 - - - 68.35 - 72.76 - - - - -
LaVIN-7B 39.32 40.75 26.84 37.22 - 33.06 - 60.48 - - - 71.41 - 69.97 - - - - -
LaVIN-13B 53.64 48.79 32.77 35.39 - 40.68 - 57.58 - - - 79.97 - 73.54 - - - - -
Lynx 64.32 67.71 42.79 46.00 - 47.00 - 43.96 - - - 74.77 - 73.72 - - - - -
Fuyu-8B 48.54 55.46 46.34 58.82 - 50.81 - 61.44 - - - 83.49 - 78.37 - - - - -
LaVIT 61.65 68.74 41.78 36.84 - 43.36 - 56.00 - - - 73.09 - 64.10 - - - - -
Qwen-VL-Chat 62.38 69.06 49.29 65.48 - 59.72 - 61.10 - - - 85.93 - 80.41 - - - - -
BLIP 66.26 68.22 51.06 70.78 - 64.42 - 72.02 - - - 88.99 - 86.88 - - - - -
InstructBLIP 71.60 70.37 52.36 57.57 59.63 60.37 35.96 73.10 - - - 88.68 - 85.98 - - - - -
LLaVA-v1.5-7B* - - - - - - - 93.67 93.40 85.18 85.11 - - - - - - - -
mPLUG-Owl2 - - - 76.80 72.30 74.20 73.00 - - - - - - - - - - - -
MMICL-14B - - - 76.00 72.70 74.10 74.00 - - - - - - - - - - - -
LLaVA-v1.5-13B - - - 77.90 74.30 74.60 74.30 - - 51.06 43.02 - - - - - - - -
Qwen-VL-v1.0 - - - - - - - - - 76.63 69.03 - - - - - - - -